OpenStudy (anonymous):

a rectangle abcd which measures 9 ft by 12ft. it folded once perpendicular to diagonal AC such that the opposite vertices A and C coincide. find the length of the fold. can someone illustrate and answer this?

OpenStudy (anonymous):

25 ft

OpenStudy (anonymous):

can you illustrate it?

OpenStudy (shinalcantara):

 By Pythagorean Theorem, the length of AC is \[(2x)^2 = 12^2 + 9^2\] \[4x^2 = 225\] \[x^2 = \frac{ 225 }{ 4 }\] \[x = 7.5\]  \[\tan \theta = \frac{ 9 }{ 12 } = \frac{ y }{ x }\] \[\frac{ 9 }{ 12 }=\frac{ y }{ 7.5 }\] \[y = \frac{ 9(7.5) }{ 12 }\] \[y = 5.625\] \[2y = 2(5.625)\] \[y = 11.25\]